OpenStudy (anonymous):

PLEASE HELP MEDAL WILL BE REWARDED!!!!!! Find an equation for the nth term of the arithmetic sequence. -17, -13, -9, -5, ... an = -17 + 4(n + 2) an = -17 x 4(n - 1) an = -17 + 4(n - 1) an = -17 + 4(n + 1) I think it might be the third one...

OpenStudy (anonymous):

that's what I think too

OpenStudy (anonymous):

common difference is 4 and the intial is -17 \[t_n=a+(n-1)d\] \[t_n=-17+(n-1)4\]

OpenStudy (anonymous):

Thank you @Chelsea04 :) What about this one?? Find an equation for the nth term of the sequence. -4, -16, -64, -256, ... an = 4 • -4n an = 4 • -4n + 1 an = -4 • 4n an = -4 • 4n - 1

OpenStudy (anonymous):

I have a feeling its the last one

OpenStudy (anonymous):

is the n-1 in brackets?

OpenStudy (anonymous):

No its like 4^n-1

OpenStudy (anonymous):

oh, then yea, the last one

OpenStudy (anonymous):

Okay thanks a lot! Im posting more questions soon if you want to earn more medals since your an amazing help!!!

OpenStudy (anonymous):

thank you :) you're so nice :)